Two cars leave towns 800 kilometers apart at the same time and travel toward each other. One car's rate is 18 kilometers per hour less than the other's. If they meet in 5 hours, what is the rate of the slower car?
Answer by rfer(16322)

800/(2x-18)=5
800=5(2x-18)
800=10x-90
10x=890
x=89 kph
x-18=71 kph